OpenStudy (anonymous):

y =-1/3x + 3 Find the slope and y-intercept.,, Medals to anyone who can help me! :)

OpenStudy (angela210793):

slope=-1/3 and y intercept is 3(just get x=0)

OpenStudy (amistre64):

y = mx + b where m = slope and b = yint

OpenStudy (anonymous):

oooh, I understand now~ Thanks guys!!!

OpenStudy (amistre64):

\(\color{green}{\text{youre welcome :)}}\)

OpenStudy (anonymous):

slope -1/3 and c=3

OpenStudy (anonymous):

its already a y intercept equation ... y=mx+c where m is the slope and c is the intercept

OpenStudy (anonymous):

How would I describe how to graph the line based on slope-intercept form?

OpenStudy (amistre64):

the slope says to go to the right by 3 and down by 1; so start at the yint (0,3) and follow the slops advice

OpenStudy (anonymous):

Thanks SOOOO MUCH!